OpenStudy (angelwings996):

Solve by using tables. Give each answer to at most two decimal places. 7x^2 + 5x - 3 = 0 a. 0.39, -1.1 b. 1.1, -0.39 c. 0.78, -2.21 d. 0.75, -0.75

OpenStudy (anonymous):

is a which is 0.39,-1.1

OpenStudy (angelwings996):

Could you tell me how you got a? So I know how to do similar problems like these

OpenStudy (anonymous):

take\[7x^2+5x-3=0\]

OpenStudy (anonymous):

\[use quadratic formula \rightarrow x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

\[a=7 \rightarrow b=5 \rightarrow c=-3\]

OpenStudy (anonymous):

put it in the formula

OpenStudy (anonymous):

this is true using this formula u will get

OpenStudy (anonymous):

\[x=(-5\pm \sqrt{25+84})/14\]

OpenStudy (anonymous):

then \[x=(-5\pm \sqrt{109})/14\]

OpenStudy (anonymous):

\[\frac{ -5pm sqrt{25-4(7)(-3)} }{ 2(7) }\]

OpenStudy (anonymous):

pm is plus minus

OpenStudy (anonymous):

\[x=(-5\pm10.4)/14\]

OpenStudy (anonymous):

\[x=0.39 or-1.1\]

OpenStudy (anonymous):

do understand ?

OpenStudy (angelwings996):

Yes I do. Thank you for your help

OpenStudy (anonymous):

dont worry !!!